# FSH3D: 3D Representation via Fibonacci Spherical Harmonics

### v1

@@ Abstract
Spherical harmonics are a favorable technique for 3D representation, employing a frequency-based approach through the spherical harmonic transform (SHT). Typically, SHT is performed using equiangular sampling grids. However, these grids are non-uniform on spherical surfaces and exhibit local anisotropy, a common limitation in existing spherical harmonic decomposition methods. This paper proposes a 3D representation method using Fibonacci Spherical Harmonics (FSH3D). We introduce a spherical Fibonacci grid (SFG), which is more uniform than equiangular grids for SHT in the frequency domain. Our method employs analytical weights for SHT on SFG, effectively assigning sampling errors to spherical harmonic degrees higher than the recovered band-limited function. This provides a novel solution for spherical harmonic transformation on non-equiangular grids. The key advantages of our FSH3D method include: 1) With the same number of sampling points, SFG captures more features without bias compared to equiangular grids; 2) The root mean square error of 32-degree spherical harmonic coefficients is reduced by approximately 34.6% for SFG compared to equiangular grids; and 3) FSH3D offers more stable frequency domain representations, especially for rotating functions. FSH3D enhances the stability of frequency domain representations under rotational transformations. Its application in 3D shape reconstruction and 3D shape classification results in more accurate and robust representations. Code will be acailable.
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010371.10010396.10010400</concept_id> <concept_desc>Computing methodologies Point-based models</concept_desc> <concept_significance>300</concept_significance> </concept> <concept> <concept_id>10002950.10003714.10003715.10003717</concept_id> <concept_desc>Mathematics of computing Computation of transforms</concept_desc> <concept_significance>300</concept_significance> </concept> </ccs2012>
@@ 1 Introduction
To address this problem, one approach is to map the model onto the faces of a suitable polyhedron, with the faces subdivided to form grid cells. Grid frameworks derived from a cube [RPM96] and those based on a regular icosahedron [MLP02] are popular. However, polyhedral subdivisions have limited ability to precisely control refinement levels. Drake $et$ $al.$ ntroduce a new method for performing spherical harmonic analysis on HEALPix data, with a computational complexity of $O(N\log^{2}N)$ [DW20]. Larkins $et$ $al.$ survey various spherical data structures for normal binning, identifying the spherical Fibonacci grid (SFG) as the most accurate method [LCD12]. While SFG provides a more uniform scheme, addressing the weight problem is crucial when applied to SHT. Swinbank $et$ $al.$ [SJP06] approximate the weight of each point on SFG as equal weights. Ahmad $et$ $al.$ [ADV07] define the area of the Voronoi diagram as the weight of each point on SFG, termed area weights. However, determining Fibonacci weights under the SHT framework poses unique challenges, requiring energy adjustment considerations for each spherical harmonic degree.
@@ 2.3 Sampling on Sphere
The challenge of evenly distributing points on a sphere, a longstanding problem, has crucial implications in numerous scientific fields. One common method for implementing spherical sampling involves mapping the model onto the faces of a polyhedron, such as mapping to a cube [RPM96] or an icosahedron [MLP02]. However, polyhedral subdivisions have limited ability to precisely control refinement levels. Another solution is spherical sampling based on a spiral pattern. HEALPix is a versatile structure for the pixelization of data on the sphere [GHB05]. SFG methods have been widely used for near-optimal solutions [NS05]. While SFG provides a more uniform scheme, addressing the weight problem is crucial when applied to SHT. Swinbank $et$ $al.$ [SJP06] approximate the weight of each point on SFG as equal weights. Ahmad $et$ $al.$ [ADV07] define the area of the Voronoi diagram as the weight of each point on SFG, termed area weights. However, determining Fibonacci weights under the SHT framework poses unique challenges, requiring energy adjustment considerations for each spherical harmonic degree.
@@ 3.2 Backgroud
where $\theta\in[0,\pi],\phi\in[0,2\pi]$ are the zenith angle and azimuth angle of the spherical coordinate system, respectively. Specifically, the spherical harmonic coefficient $\hat{f}{l}^{m}$ is computed as:
@@ 3.4 Discretization Problem of SHT
$s$ represents discrete samples on the unit sphere. The SHT is applied to $s$ to derive $\hat{s}$. The 0-degree spherical harmonic corresponds to the unit sphere, hence only the zeroth degree of $\hat{s}$ equals 1, with all other degrees being 0. As the represent process involves a band-limited function, it is essential to ensure that only finite degrees satisfy this condition. Let the cut-off degree be denoted as $e$, such that $\hat{s}\left(0,0\right)=1$ and $\hat{s}\left(l,m\right)=0$ for $0<l<e$. Thus, $s$ is expressed using spherical harmonics according to these conditions:
@@
Thus, the sampling function can be decomposed into a band-limited original function and the aliasing error introduced by sampling ${{f}{s}}$. According to Theorem 1, the product of spherical harmonics $Y_{l}^{m}Y_{j}^{k}$ can be expressed as a sum of a series of $Y_{L}^{m_{1}+m_{2}}$, where $\left|l-j\right|\leq L\leq l+j$. To effectively represent the band-limited function, the aliasing error must be confined to spherical harmonics greater than or equal to $b$, implying that $\left|l-j\right|\geq b$. With the constraints of $j$ and $l$, this condition holds true only when $e\geq 2b$. Therefore, we conclude that when at least
@@ 3.5 Determine the Weight of SHT
We employ a more uniform SFG (${{F}_{\left[\theta,\phi\right]}}$) as the sampling module for the SHT (see Sec 3.2). Typically, the SFG utilizes equal weights for calculations [SJP06]. However, each grid point of the SFG’s bounding area contains approximately 10 grid points at the extreme points, with a deviation of around 2% [SJP06]. Consequently, some researchers advocate using area weights for calculations [ADV07]. Nonetheless, when employing SFG within the SHT framework, it is essential to consider the energy of each degree of spherical harmonics. Hence, the weight cannot be directly determined by equal weights or area weights. Therefore, we utilize weighted compensation to construct the sampling function, ensuring that the sampling meets the conditions specified in Eq. (8).
@@ 4.2 Quantitative Analysis of SFG
Figure 7: Quantitative comparison of equiangular grids, SFG, HEALPix grids, and Icosahedral grids.
@@
In Fig. 7, the experiment evaluates spherical harmonic coefficients of 32 degrees across different grids. SFG, with analytic weights, exhibits the highest accuracy. Compared to equiangular grids using DH weights, SFG reduces RMSE by approximately 34.6% and MAE by approximately 34.0%. Uniformly distributed grids with equal weights address the uneven distribution of sampling points but fail to mitigate higher-degree energy errors of spherical harmonics. Consequently, grids with equal weights perform relatively poorly, yielding unstable results. The FSH3D method in this paper uses SFG with analytic weights. This approach employs a more uniform SFG, enhancing information comprehensiveness and overall stability in the SHT process.
@@ 4.3 Reconstructing Shapes with Star-shaped Sets
Figure 8: Visual comparisons of the reconstruction results of our method FSH3D (d)(f) and ESH (c)(e). The FSH3D representation excels in capturing details and remains consistent despite rotation.
@@
To assess the efficacy of FSH3D in 3D shape representation, we compile an experimental dataset. The dataset comprises 110 star-shaped shapes, manually created or sourced from the Princeton Shape Benchmark database, labeled as G1. Subsequently, all shapes in G1 are randomly rotated to produce dataset G2. In the experiment, we sample both G1 and G2 datasets into point cloud data and represent them using spherical harmonics via SHT. For comparative analysis, we employ the currently optimal equiangular grids with DH weights (ESH) alongside SFG with analytic weights (FSH3D), with both methods calculating spherical harmonic coefficients up to 32 degrees.
As demonstrated in Fig. 8, our proposed FSH3D method excels in preserving details and remains unaffected by rotation. There is a significant difference between Fig. 8(c) and Fig. 8(e). The ESH method employs uneven sampling, resulting in noticeable deformation when rotated, particularly evident as flattening at the equator in the first row. In contrast, Fig. 8(d) and Fig. 8(f) remain consistent under rotation due to FSH3D’s more uniform sampling approach. This uniformity allows FSH3D to retain shapes better compared to equiangular grids. As shown in Fig. 9, equiangular grids exhibit noticeable deformation within the red-framed area, whereas SFG provides a more accurate representation in the same region. The sparsity of the equatorial region in equiangular grids contrasts with the more uniform distribution of SFG, enabling it to capture richer details.
We use the average deviations of RMSE, MAE, and volume error (VE) for two groups of shapes before and after rotation as evaluation indicators to assess the impact of rotation on FSH3D representation. The calculation method is as follows:
@@
where, RMSE, MAE, and VE represent indicators of the deviation between the restored shape and the original shape.
@@
      $\text{RMSE}_{G1-G2}$       1.164       0.109
      $\text{MAE}_{G1-G2}$       0.486       0.049
      $\text{VE}_{G1-G2}$       1.190       0.123
For finer representation, higher degrees of spherical harmonics are necessary, which require more sampling and computation. We investigate the representation effect of SFG at higher spherical harmonic degrees. In the experiment, spherical harmonic coefficients of 32, 64, and 128 degrees are used for the shapes in G1 to perform spherical harmonic representation using both equiangular grids and SFG. We compute the numerical deviation of the spherical harmonic representation from the original shape.
@@
32 Equiangular grids 14.912 10.632 75.12
SFG 14.590 8.922 3.80
64 Equiangular grids 5.608 1.754 19.50
SFG 4.911 1.464 3.85
128 Equiangular grids 2.707 0.670 2.23
SFG 1.477 0.372 0.58

### v3

@@ Abstract
Spherical harmonics are a favorable technique for 3D representation, employing a frequency-based approach through the spherical harmonic transform (SHT). Typically, SHT is performed using equiangular sampling grids. However, these grids are non-uniform on spherical surfaces and exhibit local anisotropy, a common limitation in existing spherical harmonic decomposition methods. This paper proposes a 3D representation method using Fibonacci Spherical Harmonics (FSH3D). We introduce a spherical Fibonacci grid (SFG), which is more uniform than equiangular grids for SHT in the frequency domain. Our method employs analytical weights for SHT on SFG, effectively assigning sampling errors to spherical harmonic degrees higher than the recovered band-limited function. This provides a novel solution for spherical harmonic transformation on non-equiangular grids. The key advantages of our FSH3D method include: 1) With the same number of sampling points, SFG captures more features without bias compared to equiangular grids; 2) The root mean square error of 32-degree spherical harmonic coefficients is reduced by approximately 34.6% for SFG compared to equiangular grids; and 3) FSH3D offers more stable frequency domain representations, especially for rotating functions. FSH3D enhances the stability of frequency domain representations under rotational transformations. Its application in 3D shape reconstruction and 3D shape classification results in more accurate and robust representations. Our code is publicly available at https://github.com/Miraclelzk/Fibonacci-Spherical-Harmonics. {CCSXML} <ccs2012> <concept> <concept_id>10010147.10010371.10010396.10010400</concept_id> <concept_desc>Computing methodologies Point-based models</concept_desc> <concept_significance>300</concept_significance> </concept> <concept> <concept_id>10002950.10003714.10003715.10003717</concept_id> <concept_desc>Mathematics of computing Computation of transforms</concept_desc> <concept_significance>300</concept_significance> </concept> </ccs2012>
@@ 1 Introduction
To address this problem, one approach is to map the model onto the faces of a suitable polyhedron, with the faces subdivided to form grid cells. Grid frameworks derived from a cube [RPM96] and those based on a regular icosahedron [MLP02] are popular. However, polyhedral subdivisions have limited ability to precisely control refinement levels. Drake $et$ $al.$ introduce a new method for performing spherical harmonic analysis on HEALPix data, with a computational complexity of $O(N\log^{2}N)$ [DW20]. Larkins $et$ $al.$ survey various spherical data structures for normal binning, identifying the spherical Fibonacci grid (SFG) as the most accurate method [LCD12]. While SFG provides a more uniform scheme, addressing the weight problem is crucial when applied to SHT. Swinbank $et$ $al.$ [SJP06] approximate the weight of each point on SFG as equal weights. Ahmad $et$ $al.$ [ADV07] define the area of the Voronoi diagram as the weight of each point on SFG, termed area weights. However, determining Fibonacci weights under the SHT framework poses unique challenges, requiring energy adjustment considerations for each spherical harmonic degree.
@@ 2.3 Sampling on Sphere
The challenge of evenly distributing points on a sphere, a longstanding problem, has crucial implications in numerous scientific fields. One common method for implementing spherical sampling involves mapping the model onto the faces of a polyhedron, such as mapping to a cube [RPM96] or an icosahedron [MLP02]. However, polyhedral subdivisions have limited ability to precisely control refinement levels. Another solution is spherical sampling based on a spiral pattern. HEALPix is a versatile structure for the pixelization of data on the sphere [GHB05]. SFG methods have been widely used for near-optimal solutions [NS05]. While SFG provides a more uniform scheme, addressing the weight problem is crucial when applied to SHT. Swinbank $et$ $al.$ [SJP06] approximate the weight of each point on SFG as equal weights. Ahmad $et$ $al.$ [ADV07] define the area of the Voronoi diagram as the weight of each point on SFG, termed area weights. Reeger $et$ $al.$ [RF16] presents a novel approach to create orthogonal weights for N arbitrarily dispersed nodes by using finite differences generated based on radial basis functions. However, determining Fibonacci weights under the SHT framework poses unique challenges, requiring energy adjustment considerations for each spherical harmonic degree.
@@ 3.2 Background of spherical harmonics
where $\theta\in[0,\pi],\phi\in[0,2\pi]$ are the zenith angle and azimuth angle of the spherical coordinate system, respectively. Specifically, the spherical harmonic coefficient $\hat{f}_{l}^{m}$ is computed as:
@@ 3.4 Discretization Problem of SHT
Discrete samples on the unit sphere, denoted as $s$, undergo the SHT to obtain $\hat{s}$. The 0-degree spherical harmonic corresponds to the unit sphere, hence only the zeroth degree of $\hat{s}$ equals 1, with all other degrees being 0. As the represent process involves a band-limited function, it is essential to ensure that only finite degrees satisfy this condition. Let the cut-off degree be denoted as $e$, such that $\hat{s}\left(0,0\right)=1$ and $\hat{s}\left(l,m\right)=0$ for $0<l<e$. Thus, $s$ is expressed using spherical harmonics according to these conditions:
@@
Thus, the sampling function can be decomposed into a band-limited original function and the aliasing error introduced by sampling ${{f}_{s}}$. According to Theorem 1, the product of spherical harmonics $Y_{l}^{m}Y_{j}^{k}$ can be expressed as a sum of a series of $Y_{L}^{m_{1}+m_{2}}$, where $\left|l-j\right|\leq L\leq l+j$. To effectively represent the band-limited function, the aliasing error must be confined to spherical harmonics greater than or equal to $b$, implying that $\left|l-j\right|\geq b$. With the constraints of $j$ and $l$, this condition holds true only when $e\geq 2b$. Therefore, we conclude that when at least
@@ 3.5 Determine the Weight of SHT
We employ a more uniform SFG ${{F}_{\left[\theta,\phi\right]}}$ (see Sec 3.3) as the sampling module for the SHT (see Sec 3.2). Typically, the SFG utilizes equal weights for calculations [SJP06]. However, each grid point of the SFG’s bounding area contains approximately 10 grid points at the extreme points, with a deviation of around 2% [SJP06]. Consequently, some researchers advocate using area weights for calculations [ADV07]. Nonetheless, when employing SFG within the SHT framework, it is essential to consider the energy of each degree of spherical harmonics. Hence, the weight cannot be directly determined by equal weights or area weights. Therefore, we utilize weighted compensation to construct the sampling function, ensuring that the sampling meets the conditions specified in Eq. (8).
@@ 4.2 Quantitative Analysis of SFG
Figure 7: Quantitative comparison of equiangular grids, SFG, HEALPix grids, and Icosahedral grids. Note that the vertical scales of different charts are different, aiming to better display the details of each data.
@@
In Fig. 7, the experiment evaluates spherical harmonic coefficients of 32 degrees across different grids. SFG, with analytic weights, exhibits the highest accuracy. Compared to equiangular grids using DH weights, SFG reduces RMSE by approximately 34.6% and MAE by approximately 34.0%. Uniformly distributed grids with equal weights address the uneven distribution of sampling points but fail to mitigate higher-degree energy errors of spherical harmonics. Compared to HEALPix, SFG exhibits similar but slightly superior quality. Although Fig. 2 shows that SFG and HEALPix have similar distributions, the SFG achieves a more uniform distribution over the entire sphere. Consequently, grids with equal weights perform relatively poorly, yielding unstable results. The FSH3D method in this paper uses SFG with analytic weights. This approach employs a more uniform SFG, enhancing information comprehensiveness and overall stability in the SHT process.
@@ 4.3 Reconstructing Shapes with Star-shaped Sets
Figure 8: Visual comparisons of the reconstruction results of our method FSH3D (d)(f) and ESH (c)(e). (a) represents the input point cloud data, while (b) shows the corresponding mesh used as the ground truth. (c) and (d) depict the reconstruction results using (a) as the input data, whereas (e) and (f) present the reconstruction results when (a) is rotated by $90^{\circ}$ and used as the input. Deviations are measured using the Hausdorff distance, with the color map indicating reconstruction quality.
@@
To assess the efficacy of FSH3D in 3D shape representation, we compile an experimental dataset. The dataset comprises 110 star-shaped shapes, manually created or sourced from the Princeton Shape Benchmark database, labeled as G1. We normalize each shape in G1 according to the size of the bounding box, and the dataset G2 is generated by randomly rotating the shapes in G1. This is used to compare the effect of random rotation on the performance of FSH3D. In the experiment, we sample both G1 and G2 datasets into point cloud data and represent them using spherical harmonics via SHT. For comparative analysis, we employ the currently optimal equiangular grids with DH weights (ESH) alongside SFG with analytic weights (FSH3D), with both methods calculating spherical harmonic coefficients up to 32 degrees.
We use the average deviations of RMSE, MAE, and volume error (VE) for two groups (G1 and G2) of shapes before and after rotation as evaluation indicators to assess the impact of rotation on FSH3D representation. The calculation method is as follows:
@@
where, RMSE and MAE for each group of shapes are calculated by statistically analyzing the deviation in Hausdorff distance between the restored and original shapes. The VE is determined by calculating the volume deviation between the restored and original shapes.
@@
      $\text{RMSE}_{G1-G2}$       1.083       0.087
      $\text{MAE}_{G1-G2}$       0.443       0.038
      $\text{VE}_{G1-G2}$       1.072       0.114
Fig. 8 shows the visual results of reconstructions of our method FSH3D, with all shapes are normalized and visualized from the same perspective. Our proposed method FSH3D excels in preserving details and remains unaffected by rotation. There is a significant difference between Fig. 8(c) and Fig. 8(e). The ESH method employs uneven sampling, resulting in noticeable deformation when rotated, particularly evident as flattening at the equator in the first row. In contrast, Fig. 8(d) and Fig. 8(f) remain consistent under rotation due to FSH3D’s more uniform sampling approach. This uniformity allows FSH3D to retain shapes better compared to equiangular grids. As shown in Fig. 9, equiangular grids exhibit noticeable deformation within the red-framed area, whereas SFG provides a more accurate representation in the same region. The sparsity of the equatorial region in equiangular grids contrasts with the more uniform distribution of SFG, enabling it to capture richer details.
For finer representation, higher degrees of spherical harmonics are necessary, which require more sampling and computation. We investigate the representation effect of SFG at higher spherical harmonic degrees, using G1 as the experimental data. In the experiment, spherical harmonic coefficients of 32, 64, and 128 degrees are used for the shapes in G1 to perform spherical harmonic representation using both equiangular grids and SFG. We compute the numerical deviation of the spherical harmonic representation from the original shape.
@@
32 Equiangular grids $14.902$ $10.612$ $75.12$
SFG $14.577$ $8.901$ $3.80$
64 Equiangular grids $5.598$ $1.754$ $19.50$
SFG $4.886$ $1.457$ $3.85$
128 Equiangular grids $2.702$ $0.665$ $2.23$
SFG $1.467$ $0.353$ $0.58$
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 92367301, No. 92267201, No. 52275493, No. 92160301).